I have 500g of water at 25oC. If the amount of heat required to raise the temperature of this water to 90oC is the equivalent amount of kinetic energy that a 86-kg ball possesses, what is its speed?

options:

-89.4 m/s.

-77.7 m/s

-56.2 m/s

-32.8 m/s
Answer #1

Heat Q = msat - 500 x 4 186 x 65 m=500gm S= 4.186 51m BT = (90-250°C = 65c = 136045 I & DRE (1360455 = Y mae 136045² x 86x6 [